# Generative Kaleidoscopic Networks

### v1

@@ Abstract
We discovered that the Deep ReLU networks (or Multilayer Perceptron architecture) demonstrate an ‘over-generalization’ phenomenon. That is, the output values for the inputs that were not seen during training are mapped close to the output range that were observed during the learning process. In other words, the MLP learns a many-to-one mapping and this effect is more prominent as we increase the number of layers or depth of the MLP. We utilize this property of Deep ReLU networks to design a dataset kaleidoscope, termed as ‘Generative Kaleidoscopic Networks’. Briefly, if we learn a MLP to map from input $x\in\mathbb{R}^{D}$ to itself $f_{\mathcal{N}}(x)\rightarrow x$, the Kaleidoscopic sampling procedure starts with a random input noise $z\in\mathbb{R}^{D}$ and recursively applies $f_{\mathcal{N}}(\cdots f_{\mathcal{N}}(z)\cdots)$. After a burn-in period duration, we start observing samples from the input distribution and we found that deeper the MLP, higher is the quality of samples recovered. Scope: We observed this phenomenon to various degrees for the other deep learning architectures like CNNs, Transformers & U-Nets and we are currently investigating them further.
@@ 1 Introduction
Figure 1: Manifold learning & Kaleidoscopic sampling. [left] During the manifold learning process, the Multilayer Perceptron (MLP) weights have their gradients enabled, shaded in blue, have their output units bounded between $(0,1)$ via a ‘Sigmoid’ non-linearity (or between $(-1,1)$ for ‘Tanh’), and are learned as per Eq. 1. [right] During the sampling process, the weights of the neural network model are frozen and the input is a randomly initialized in the input space, e.g. it can be sampled from a normal distribution. The model $f$ is repeatedly applied to the input in accordance to Eq. 2. Once the function is applied ‘B’ number of times $f\circ f\circ\cdots\circ f$, we start obtaining the samples closer to the input data distribution. The samples obtained after the burn-in period, $B\rightarrow K$, are the valid samples obtained from the model.
@@ 2.1 Extrapolating neural network’s behavior in higher-dimensions
Validating the ‘dimple manifold’ concept. We digress a bit and address an important observation made in this work [7] about NNs learning dimple manifolds. Based on the loss function vs number of points curve over different dimensions of input, refer Fig. 6, we can experimentally say that the dimple manifold proposal that the NNs hyperplane creates a ‘dimple’ (or a dip, or a small trench) around the training data points. Hence, with slight change in the NN weights, one can change the NN output for a given input data point. As the loss function curve is relatively flat, there is not much loss difference between the observed (training) and the unobserved (testing) data points. This also validates the general observation that one can adversarily attack/trick the neural network by slightly perturbing the input. From the Fig. 2,3&5, we can see that a smooth manifold dip is formed on the loss manifold at the training points. Now, if there are many such training points, then the plots in Fig. 6 suggests that the manifold will have a very low loss values throughout the space. Thus, when adversarily attacking the model, one can make slight modifications to the input to increase its loss value and thereby the point is pulled out from the manifold dip and gets misclassified.
@@ 3 Generative Kaleidoscopic  Networks
Observations: We revisit the Fig. 2,3&5 which analyze points in 1D and 2D spaces. We observe from the loss manifold profiles that one can get also get multiple minima in the regions where the training data is not present. This means that the probability of getting samples from that region becomes high, which is not desired. When we see the corresponding output manifold profiles and sampling results, we can see that if we repeatedly apply the model $f_{\mathcal{N}}$, one can get samples from the training data distribution but will require more number of iterations (or longer burn-in periods). The Kaleidoscopic sampling procedure will improve if we can make the output manifold look like a step-function over the points observed in training distributions which happens when we increase the depth of the MLP. Fig. 7 shows the Kaleidoscopic sampling results for MNIST images. In our mind, we visualize that the output vs input profile of the MLP as multiple steps (many-to-one mappings), with each step containing the distribution of a digit. Thus, we can start with any input random noise and end up on one of these digit steps after the burn-in period of the Kaleidoscopic sampling.
@@
Are we just memorizing? The deeper the MLP, the stronger is the many-to-one mapping, which in turn makes the output vs input manifold look more like a step function. Since, MLP is not able to fit well (reduce manifold learning loss) over the image data as compared to other architectures like CNNs & Transformers, the quality of samples obtained are not that good. But, for MNIST images, we do get decent loss convergence and we visually analyse the samples recovered and also measured distance from the closest training images and our initial conclusion suggests that it is not just purely memorizing. We added ‘MNIST Kaleidoscope’ visualizations in the software link.
@@ 4 Related Works
Gradient based sampling: We consider the input to be a learnable tensor and then use a gradient descent based sampling procedure. Variants of this approach have been used previously as well in other contexts for doing adversarial attacks [12] or learning and sampling probabilistic graphical models [8]. The idea is to fix the learned model weights and tune the input tensor to match the model output. Essentially, we optimize for the following loss $\operatorname*{arg\,min}_{\hat{X}}\frac{1}{D}\left\|\hat{X}-f_{\mathcal{N}}(\hat{X})\right\|_{2}^{2}$ where, the input learnable tensor $\hat{X}$ is initialize randomly, for instance, from a Gaussian distribution $\sim\mathcal{N}(0,I)$. It is a parallel approach to Kaleidoscopic sampling, which is based on learning the input tensor via gradient descent while the model is frozen. It seemed to work well on an Infant Mortality data [11] for Neural Graphical Models [8] and their follow up work on Federated learning [2]. We also want to underscore that the choice of architecture plays an important role in the quality of samples generated. We tried these gradient-based sampling approaches for image generation using Multilayer Perceptron, Convolutional Neural Networks & Transformer based architectures. In general, we found that one will not always obtain a sample that looks similar to the input training data even though loss value goes down considerably. In other words, while running the sampling optimization, the input learnable tensor and model output matches, but the learned tensor does not visually match the training images (sometimes looks like noise). This suggests that the manifold learned by fitting the model to the input data points is too vast. Instead of randomly initializing $\hat{X}$, if we add some white noise to a data point from the training data, the sampling technique suggested, seems to work better.
@@
Figure 7: Generative Kaleidoscopic Network for MNIST: We do manifold learning on a MLP with $L=10,H=500$ with intermediate ‘ReLU’ and final layer with ‘Tanh’ non-linearity. We create a ‘MNIST Kaleidoscope’ by setting $\epsilon=0.01$ in the Kaleidoscopic sampling procedure as described in Eq. 3. [TOP row] The leftmost images show the digits recovered after the manifold learning. The center images show the $1^{st}$ run after applying learned MLP on the input noise. The input noise vector was sampled randomly from a Uniform distribution ${}\mathcal{U}(-1,1)$. (Note, we get similar results with Normal distribution too). The rightmost images show the state at the sampling run of 300. We note that the procedure can converge at a digit and then it can remain stable throughout the future iterations as it has found a stable minima or step as defined in our analysis. For this reason, we add a slight noise at every sampling iteration to simulate kaleidoscopic behaviour. Here, we show the output of randomly chosen subset of images and the seed is constant, so that there is one-to-one correspondence across sampling runs. [MIDDLE row] A sequence of intermediate sampling runs $5\rightarrow 13$, which is still in the burn-in period. One can take a digit and observe their evolution over the iterations. For instance, top row & third last column evolves into digit $8$ and then ultimately becomes digit $6$ at the $300^{th}$ iteration. [BOTTOM row] Shows another sequence of evolution of sevens from noise observed.
@@ 5 Conclusions and Future Work
We observed an ‘over-generalization’ phenomenon in Neural Networks. We utilized it to introduce a new paradigm of generative modeling by creating a dataset specific ‘Kaleidoscope’. Akin to fractal generation, we start with an input noise and our Kaleidoscopic sampling procedure generates samples by doing recursive functional calls. At the current stage, without any post-processing, we emphasize that our sampling results are not close to state-of-the-art Generative AI models. This can change, if we can better understand the manifold profiles for more specialized architectures used for image/audio/video analysis. We are exploring theoretical explanations, experiments on tabular data, as well as conditional generation using the Generative Kaleidoscopic Networks.

### v4

@@ Abstract
We discovered that the neural networks, especially the deep ReLU networks, demonstrate an ‘over-generalization’ phenomenon. That is, the output values for the inputs that were not seen during training are mapped close to the output range that were observed during the learning process. In other words, the neural networks learn a many-to-one mapping and this effect is more prominent as we increase the number of layers or depth of the neural network. We utilize this property of neural networks to design a dataset kaleidoscope, termed as ‘Generative Kaleidoscopic Networks’. Briefly, if we learn a model to map from input $x\in\mathbb{R}^{D}$ to itself $f_{\mathcal{N}}(x)\rightarrow x$, the proposed Kaleidoscopic sampling procedure starts with a random input noise $z\in\mathbb{R}^{D}$ and recursively applies $f_{\mathcal{N}}(\cdots f_{\mathcal{N}}(z)\cdots)$. After a burn-in period duration, we start observing samples from the input distribution and the quality of samples recovered improves as we increase the depth of the model. Scope: We observed this phenomenon to various degrees for the other deep learning architectures like CNNs, Transformers & U-Nets and we are currently investigating them further.
@@ 1 Introduction
Figure 1: Manifold learning & Kaleidoscopic sampling. [left] During the manifold learning process, the deep ReLU networks or the Multilayer Perceptron (MLP) weights have their gradients enabled, shaded in blue, and their output units bounded between $(0,1)$ via a ‘Sigmoid’ non-linearity (or between $(-1,1)$ for ‘Tanh’). They are learned as per Eq. 1. [right] During the sampling process, the weights of the neural network model are frozen and the input is a randomly initialized noise, sampled from a normal or uniform distribution. The model $f$ is repeatedly applied to the input noise in accordance to Eq. 2. Once the function is applied ‘B’ number of times $f\circ f\circ\cdots\circ f$, we start obtaining the samples closer to the input data distribution $B\rightarrow K$.
@@ 2.1 Extrapolating neural network’s behavior in higher-dimensions
Validating the ‘dimple manifold’ concept. We digress a bit and address an important observation made in this work [12] about NNs learning dimple manifolds. Based on the loss function vs number of points curve over different dimensions of input, refer Fig. 6, we can experimentally say that the dimple manifold proposal that the NNs hyperplane creates a ‘dimple’ (or a dip, or a small trench) around the training data points. Hence, with slight change in the NN weights, one can change the NN output for a given input data point. As the loss function curve is relatively flat, there is not much loss difference between the observed (training) and the unobserved (testing) data points. This also validates the general observation that one can adversarily attack/trick the neural network by slightly perturbing the input. From the Fig. 2,3,4&5, we can see that a smooth manifold dip is formed on the loss manifold at the training points. Now, if there are many such training points, then the plots in Fig. 6 suggests that the manifold will have a very low loss values throughout the space. Thus, when adversarily attacking the model, one can make slight modifications to the input to increase its loss value and thereby the point is pulled out from the manifold dip and gets misclassified.
@@ 3 Generative Kaleidoscopic  Networks
Observations: We revisit the Fig. 2,3,4&5 which analyze points in 1D and 2D spaces. We observe from the loss manifold profiles that one can get also get multiple minima in the regions where the training data is not present. This means that the probability of getting samples from that region becomes high, which is not desired. When we see the corresponding output manifold profiles and sampling results, we can see that if we repeatedly apply the model $f_{\mathcal{N}}$, one can get samples from the training data distribution but will require more number of iterations (or longer burn-in periods). The Kaleidoscopic sampling procedure will improve if we can make the output manifold look like a step-function over the points observed in training distributions which happens when we increase the depth of the MLP. Fig. 7 shows the Kaleidoscopic sampling results for MNIST images. In our mind, we visualize that the output vs input profile of the MLP as multiple steps (many-to-one mappings), with each step containing the distribution of a digit. Thus, we can start with any input random noise and end up on one of these digit steps after the burn-in period of the Kaleidoscopic sampling.
@@
Are we just memorizing? Depends on the Dataset Manifold. The deeper the MLP, the stronger is the many-to-one mapping, which in turn makes the output vs input manifold look more like a step function. For MNIST images, refer Fig. 7, we get decent loss convergence and we visually can observe the ‘smooth’ transitions among the samples recovered. We further measured distance from the closest training images and our initial conclusion suggests that it is not just purely memorizing. We further ran experiments on a subset of CIFAR and CELEB-A datasets, refer Fig. 8. In some cases, we can observe some ‘jumpy’ transitions between the Kaleidoscopic sampling iterations. Our initial assessment to explain this observations is that more steps in the neural network manifolds are formed if the dataset contains disconnected points and this indicates that the dataset is a collection from distinct distributions. In some sense, we can detect whether the dataset lies in a smooth manifold and whether there is ‘semantic’ similarity among the datapoints. Another interesting aspect to explore is doing data compression using these networks. In order to help reader understand the sample transition reasoning discussed here, we added various ‘Dataset Kaleidoscopes’ animations in the software link.
@@ 4 Related Works
Gradient based sampling: We consider the input to be a learnable tensor and then use a gradient descent based sampling procedure. Variants of this approach have been used previously as well in other contexts for doing adversarial attacks [25] or learning and sampling probabilistic graphical models [20, 19, 18, 21]. The idea is to fix the learned model weights and tune the input tensor to match the model output. Essentially, we optimize for the following loss $\operatorname*{arg\,min}_{\hat{X}}\frac{1}{D}\left\|\hat{X}-f_{\mathcal{N}}(\hat{X})\right\|_{2}^{2}$ where, the input learnable tensor $\hat{X}$ is initialize randomly, for instance, from a Gaussian distribution $\sim\mathcal{N}(0,I)$. It is a parallel approach to Kaleidoscopic sampling, which is based on learning the input tensor via gradient descent while the model is frozen. It seemed to work well on an Infant Mortality data [24] for Neural Graphical Models [20] and their follow up work on Federated learning [3]. We also want to underscore that the choice of architecture plays an important role in the quality of samples generated. We tried these gradient-based sampling approaches for image generation using Multilayer Perceptron, Convolutional Neural Networks & Transformer based architectures. In general, we found that one will not always obtain a sample that looks similar to the input training data even though loss value goes down considerably. In other words, while running the sampling optimization, the input learnable tensor and model output matches, but the learned tensor does not visually match the training images (sometimes looks like noise). This suggests that the manifold learned by fitting the model to the input data points is too vast. Instead of randomly initializing $\hat{X}$, if we add some white noise to a data point from the training data, the sampling technique suggested, seems to work better.
@@
Figure 7: Generative Kaleidoscopic Network for MNIST: We do manifold learning on a MLP with $L=10,H=500$ with intermediate ‘ReLU’ and final layer with ‘Tanh’ non-linearity. We create a ‘MNIST Kaleidoscope’ by setting $\epsilon=0.01$ in the Kaleidoscopic sampling procedure as described in Eq. 3. [TOP row] The leftmost images show the digits recovered after the manifold learning. The center images show the $1^{st}$ run after applying learned MLP on the input noise. The input noise vector was sampled randomly from a Uniform distribution ${}\mathcal{U}(-1,1)$. (Note, we get similar results with Normal distribution too). The rightmost images show the state at the sampling run of 300. We note that , at times, the procedure can converge at a digit and then it can remain stable throughout the future iterations as it has found a stable minima or step as defined in our analysis. For this reason, we add a slight noise at every sampling iteration to simulate kaleidoscopic effect. Here, we show the output of randomly chosen subset of images and the seed is constant, so that there is one-to-one correspondence across sampling runs. [MIDDLE row] A sequence of intermediate sampling runs $5\rightarrow 13$, which is still in the burn-in period. One can take a digit and observe their evolution over the iterations. For instance, top row & third last column evolves into digit $8$ and then eventually transitions to digit $6$ at the $300^{th}$ iteration. [BOTTOM row] Shows another sequence of evolution of sevens from noise.
Figure 8: Generative Kaleidoscopic Networks: CIFAR & CELEB-A. We create dataset kaleidoscopes on a subset of the images. We do manifold learning with MLP $L=10,H=2000$. [TOP] We randomly chose 1K dog images from CIFAR-10 dataset. The leftmost image shows some images from the training data, while the remaining images are the samples obtained during the intermediate sampling transitions starting with initial random noise. [BOTTOM] We randomly chose 1K celebrity images from the Celeb-A dataset. In this case, we observe more ‘jumpy’ transitions between sampling and the convergence in term of the sampling iterations also occurs quicker. Our initial reasoning for this is that the image manifold is not smooth and hence there are more steps in the output manifold causing the sampling procedure to jump abruptly or get stuck at a sample. (Animations included in the demo link)
@@ 5 Conclusions and Future Work
We observed an ‘over-generalization’ phenomenon in neural networks. We utilized it to introduce a new paradigm of generative modeling by creating a dataset specific ‘Kaleidoscope’. Akin to fractal generation, we start with an input noise and our Kaleidoscopic sampling procedure generates samples by doing recursive functional calls. At the current stage, without any post-processing, we emphasize that our sampling results do not exceed the state-of-the-art Generative AI models. This can change, if we can better understand the manifold profiles for more specialized architectures used for image/audio/video analysis. We are exploring theoretical explanations, experiments on multimodal data, as well as conditional generation using the Generative Kaleidoscopic Networks.
Applications in video processing: Video-based tasks such as action recognition [10], video segmentation [26], object detection [7], identifying activity within videos [1] and others are of great interest to the research community and have a considerable commercial impact [9]. A natural application of the over-generalization phenomenon in conjunction with Kaleidoscopic sampling seems to be in the video representation and generation domain. Videos can be seen as a sequential collection of frames that are closely related. Our initial experimentation shows that the frame embeddings of a given video lie very close to each other in the manifold space. We are currently investigating Generative Kaleidoscopic Network’s compatibility with the video prior representation frameworks provided in [16, 17]. The Kaleidoscopic sampling procedure can be thought of as a stochastic process (or a temporal process). This idea can be extended to represent continuous multi-dimensional processes as done for videos in [15]. Furthermore, we are investigating diverse video generation tasks and the techniques provided in the following works [4, 14, 13] seem to be good initial candidates.